# GenAD: Generative End-to-End Autonomous Driving

### 20 Feb 2024 v2

@@ 1 Introduction
Vision-centric autonomous driving has been extensively explored in recent years due to its economic convenience [42, 24, 39, 16, 25]. While researchers have advanced the limit of vision-centric autonomous driving in various tasks including 3D object detection [24, 16, 25], map segmentation [28, 32, 23, 52], and 3D semantic occupancy prediction [17, 49, 44, 45, 48, 55], recent advances in vision-centric end-to-end autonomous driving [15, 14, 19] have shed light on a potential and elegant path to directly produce planning results from raw sensors.
@@
Most existing end-to-end autonomous driving models are composed of several modules and follow a pipeline of perception, motion prediction, and planning [15, 14, 19]. For example, UniAD [15] further progressively performs map perception, detection, tracking, motion prediction, occupancy prediction, and planning modules to improve the robustness of the system. It is also observed that using a planning objective improves the performance of intermediate tasks [15]. However, the serial design of prediction and planning of existing pipelines ignores the possible future interactions between the ego car and the other traffic participants. We argue that this type of interaction is important for accurate planning. For example, the lane shift of the ego car would affect the action of the rear cars, and further affects the planning of the ego car. This high-order interaction cannot be effectively modeled by the current design of motion prediction before planning. Also, future trajectories are highly structured and share a common prior (e.g., most trajectories and continuous and straight lines). Still, most existing methods fail to consider this structural prior, leading to inaccurate predictions and planning.
@@ 2 Related Work
Perception. Perception is the basic step in autonomous driving, which aims to extract meaningful information from raw sensor inputs. Despite the strong performance of LiDAR-based methods [53, 4, 36, 8], vision-centric methods [17, 51, 49, 24, 25] have emerged as a competitive alternative due to the low costs of RGB cameras. Equipped with large 2D image backbones, vision-centric methods demonstrated great potential in the main perception tasks including 3D object detection [42, 24, 39, 16, 51, 25], HD map reconstruction [28, 32, 23, 52], and 3D semantic occupancy prediction [49, 44, 45, 48, 55]. To accurately complete these 3D tasks, the key procedure is to transform image features to the 3D space. One line of works predicts explicit depths for image features and then projects them into the 3D space using camera parameters [42, 24, 33, 27, 39, 16, 51]. Other methods initialize queries in the 3D space and exploit deformable cross-attention to adaptively aggregate information from 2D images [25, 20, 17]. Some works further design better positional embedding strategies [31], 3D representations [17], or task heads [51] to further improve the perception performance or efficiency. In this paper, we adopt conventional simple designs for 3D perception and focus on motion prediction and planning.
@@
Planning. Planning is the ultimate goal for the first stage of autonomous driving. Despite the mature development of rule-based planners [46, 1, 7], learning-based planners [43, 5, 40] are receiving increasing attention due to their great potential to benefit from large-scale driving data and compatibility with end-to-end autonomous driving methods. Most existing end-to-end planning methods follow a pipeline of perception, prediction, and planning [15, 14, 19]. For example, ST-P3 [14] progressively employs a map perception, a BEV occupancy prediction, and a trajectory planning module to obtain future ego movements from surrounding cameras. UniAD [15] further extends ST-P3 with additional detection, tracking, and motion prediction modules to improve the robustness of the system. VAD [19] simplifies UniAD with vectorized scene representation and only map, motion, and planning modules for end-to-end driving, which achieves state-of-the-art planning performance with better efficiency. However, the serial design of prediction and planning ignores the effect of future ego movements on the agent motion prediction. It also lacks modeling of the uncertain nature of motion prediction and planning. To address this, GenAD models autonomous driving in a generative framework and simultaneously generates the future trajectories for the ego vehicle and other agents in a learned probabilistic latent space.

### 07 Apr 2024 v3

@@ 1 Introduction
Vision-centric autonomous driving has been extensively explored in recent years due to its economic convenience [44, 26, 41, 17, 27]. While researchers have advanced the limit of vision-centric autonomous driving in various tasks including 3D object detection [26, 17, 27], map segmentation [30, 34, 25, 53], and 3D semantic occupancy prediction [19, 51, 46, 47, 50, 57, 18], recent advances in vision-centric end-to-end autonomous driving [16, 15, 21] have shed light on a potential and elegant path to directly produce planning results from raw sensors.
@@
Most existing end-to-end autonomous driving models are composed of several modules and follow a pipeline of perception, motion prediction, and planning [16, 15, 21, 5]. For example, UniAD [16] further progressively performs map perception, detection, tracking, motion prediction, occupancy prediction, and planning modules to improve the robustness of the system. It is also observed that using a planning objective improves the performance of intermediate tasks [16]. However, the serial design of prediction and planning of existing pipelines ignores the possible future interactions between the ego car and the other traffic participants. We argue that this type of interaction is important for accurate planning. For example, the lane shift of the ego car would affect the action of the rear cars, and further affects the planning of the ego car. This high-order interaction cannot be effectively modeled by the current design of motion prediction before planning. Also, future trajectories are highly structured and share a common prior (e.g., most trajectories are continuous and straight lines). Still, most existing methods fail to consider this structural prior, leading to inaccurate predictions and planning.
@@ 2 Related Work
Perception. Perception is the basic step in autonomous driving, which aims to extract meaningful information from raw sensor inputs. Despite the strong performance of LiDAR-based methods [55, 4, 38, 9], vision-centric methods [19, 53, 51, 26, 27] have emerged as a competitive alternative due to the low costs of RGB cameras. Equipped with large 2D image backbones, vision-centric methods demonstrated great potential in the main perception tasks including 3D object detection [44, 26, 41, 17, 53, 27], HD map reconstruction [30, 34, 25, 53], and 3D semantic occupancy prediction [51, 46, 47, 50, 57, 54]. To accurately complete these 3D tasks, the key procedure is to transform image features to the 3D space. One line of works predicts explicit depths for image features and then projects them into the 3D space using camera parameters [44, 26, 35, 29, 41, 17, 53]. Other methods initialize queries in the 3D space and exploit deformable cross-attention to adaptively aggregate information from 2D images [27, 22, 19]. Some works further design better positional embedding strategies [33], 3D representations [19], or task heads [53] to further improve the perception performance or efficiency. In this paper, we adopt conventional simple designs for 3D perception and focus on motion prediction and planning.
@@
Planning. Planning is the ultimate goal for the first stage of autonomous driving. Despite the mature development of rule-based planners [48, 1, 8], learning-based planners [45, 6, 42] are receiving increasing attention due to their great potential to benefit from large-scale driving data and compatibility with end-to-end autonomous driving methods. Most existing end-to-end planning methods follow a pipeline of perception, prediction, and planning [16, 15, 21, 5]. For example, ST-P3 [15] progressively employs a map perception, a BEV occupancy prediction, and a trajectory planning module to obtain future ego movements from surrounding cameras. UniAD [16] further extends ST-P3 with additional detection, tracking, and motion prediction modules to improve the robustness of the system. VAD [21] simplifies UniAD with vectorized scene representation and only map, motion, and planning modules for end-to-end driving, which achieves state-of-the-art planning performance with better efficiency. However, the serial design of prediction and planning ignores the effect of future ego movements on the agent motion prediction. It also lacks modeling of the uncertain nature of motion prediction and planning. To address this, GenAD models autonomous driving in a generative framework and simultaneously generates the future trajectories for the ego vehicle and other agents in a learned probabilistic latent space.